Given the obtained statistic is t = + 2.7 alpha a = 0.05 , 1 tail ( sample>pop) . There were 17 subjects. What is the correct answr? Retain null hypothesis? Retain alternative hypothesis? Reject null hypothesis? reject alternative hypothesis?
Answer #1

There were 17 subjects ie n=17 Given, obtained statistie t= 2.57 with degrees of freedom (n-1) =(17-1)=16 It is a One failedWe reject the null hypothesis.